OpenStudy (moss65):

consider the set A=[0,1)U(1,2]. 1) Find a finite open cover of A. 2) Show that A is not compact by using Heine-Borel theorem. 3) show that A is not compact by finding an open cover of A with no finite subcover.

OpenStudy (anonymous):

1) \(\{\mathbb{R}\}\) certainly covers \(A\) and is a finite open cover.

OpenStudy (anonymous):

Do you remember the Heine-Borel theorem?

OpenStudy (anonymous):

It states that in \(\mathbb{R}\) a set is closed and bounded if and only if it is compact.

OpenStudy (anonymous):

In other words, closed and boundedness is equivalent to compactness in \(\mathbb{R}\).

OpenStudy (anonymous):

So what does that tell us about 2?

OpenStudy (moss65):

Yes, I do. #2 is not problem because it's obvious. the problem is #3

OpenStudy (anonymous):

Ok

OpenStudy (moss65):

Is {R} finite or infinite?

OpenStudy (anonymous):

A finite cover or finite subcover refers to the number of elements in the cover.

OpenStudy (anonymous):

Let's discuss 3 and you will see what I mean.

OpenStudy (anonymous):

The idea is to take a sequence of open sets that approaches 1. That way it is impossible to include only a finite number of sets in your subcover since you will not cover everything if you do.

OpenStudy (anonymous):

\(\mathscr{C} = \{(-\infty, 1 - \frac{1}{n}), (1 + \frac{1}{n}, \infty) : n \in \mathbb{N}\}\)

OpenStudy (moss65):

I see. Could it be the union of intervals let say (-1,1-1/n)U(1+1/n, 3)?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

You understand why it works?

OpenStudy (anonymous):

The union of the elements of the cover certainly contains \(A\) so it is a valid cover.

OpenStudy (anonymous):

However, it is impossible to take only a finite number of the elements to cover \(A\).

OpenStudy (moss65):

it works because intervals are close to 1 but don't include it. Am I right?

OpenStudy (moss65):

why not? we need to find finite open cover

OpenStudy (anonymous):

Compactness states that every open cover has a finite subcover.

OpenStudy (anonymous):

We have just found an open cover which does not have any finite subcover. Do you understand why there is no finite subcover?

OpenStudy (moss65):

no yet

OpenStudy (anonymous):

Can you find a finite subset of \(\mathscr{C}\) which covers \(A\)? If not, why?

OpenStudy (moss65):

Wait. we need to show that A is not compact by finding open cover with no finite subcovers. if there are finite subcovers, the set is compact. Yes?

OpenStudy (anonymous):

Compactness states that every open cover has a finite subcover. So to prove that a set is compact you must show that every possible open cover has a finite subcover. On the other hand to show a set is NOT compact, you must only provide one open cover that has no finite subcover.

OpenStudy (anonymous):

Once again we return to our solution: \(\mathscr{C} = \{(-\infty, 1 - \frac{1}{n}), (1 + \frac{1}{n}, \infty) : n \in \mathbb{N}\}\) We claim it is an open cover, but it has no finite subcover. If this is true then \(A\) is not compact.

OpenStudy (anonymous):

Let is carefully define what the terms mean to avoid confusion. An open cover for some set \(A\) is a collection of open sets, \(\mathscr{C}\), whose union contains \(A\). A finite subcover of \(\mathscr{C}\) is a finite subset \(C \subset \mathscr{C}\) such that \(C\) is still a cover for \(A\). In other words it is a finite subset of \(\mathscr{C}\) whose union contains \(A\). Do you understand the definitions of these terms?

OpenStudy (moss65):

Yes, but we need to find in #1 open finite cover. i guess it is some open interval which differs from (-inf, inf)

OpenStudy (anonymous):

1) Is a bit strange and rather trivial. You usually don't talk about finite open covers. An open cover of some set \(A\) is simply any collection of open sets whose union contains \(A\). A finite open cover is simply a finite collection meeting the above requirements.

OpenStudy (anonymous):

So let me ask you, is \(\mathscr{C} = \{\mathbb{R}\}\) an open cover for \(A\)?

OpenStudy (moss65):

it's open, but not finite

OpenStudy (anonymous):

Finite just means the cover has a finite number of elements. Tell me, how many elements does \(\mathscr{C}\) have?

OpenStudy (moss65):

Do you mean just one?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

If you are troubled by including the entire space, another solution would be \(\mathscr{C} = \{(-1, 3)\}\). But both are valid solutions.

OpenStudy (anonymous):

In any case, I think question #1 is simply there to test your understanding of the definitions. It is trivial.

OpenStudy (anonymous):

Do you understand the solution for 3?

OpenStudy (moss65):

I think I do. I just need to reread some theoretical stuff. Thank you so much for help and clarifying.

OpenStudy (anonymous):

If you have any questions about definitions, you can certainly continue to ask here. Try to carefully understand what compactness means. To understand compactness you also need to understand what an open cover is and what a finite subcover is.